two cars traveling 725mi apart and traveling toward each other along the same road. they meet in 5 hours. one car is traveling 7mph slower than the other. what is the speed of each car?
Answer by Boreal(15235)

d=r*t, t=5 hours
x mph is speed of one car/hr, x-7 mph the speed of the other/hr,
in 5 hours, their combined distance will be 725 miles
so 5(x)+5(x-7)=725
5x+5x-35=725
10x=760
x=76 mph, x-7=69 mph.
In 5 hours, one car will travel 380 miles, the other 345 miles, which add to 725 miles